Ticket Number: HEA-NOC/20070228-2 Ticket Status: OPEN
Ticket Type: unscheduled Resolver:
Ticket Opened: 20070228 09:27 UTC Problem Start: 20070227 19:59 UTC
Ticket Update: 20070228 09:27 UTC Problem End:
NUIM's primary link went down at approx 19:59 last night. Flap seen on
other BT ATM circuits at the same time. NUIM not aware of any local
activity. HEAnet not aware of any planned activity in BT. NUIM backup link
is currently ISDN, so connectivity is degraded.
20070228 09:16 BS spoke to NUIM, opening ticket with BT.
This ticket can be monitored at http://www.hea.net/tickets/20070228-2